# Design and Performance of a Universal SiPM Readout System for X- and Gamma-Ray Missions

### v1

@@ Abstract
Gamma-rays, X-rays, Spectrometry, Scintillators, Polarization, Space Qualification
@@ 1 Introduction
The joint Gravitational Wave (GW) and Gamma-ray Burst (GRB) detection of GW170817/GRB170817A indicated that gamma-ray detectors are vital in the field of multi-messenger astrophysics. These detectors need to be capable of detecting weak transients and alerting the community with temporal, spectral and localization information. This increase in interest in developing such detectors coincided with the emergence of SiPM technology. Traditional gamma-ray detectors, such as BATSE (Fishman et al., 1994), Fermi-GBM (Meegan et al., 2009), and POLAR (Produit et al., 2018), were based on scintillators read out by Photo-Multiplier Tubes (PMT). Downsides of PMTs are that they are bulky and mechanically fragile. In addition, such instruments required complex high-voltage power supplies capable of delivering the typical kV level voltages. The development of SiPMs mitigates all such issues as they are compact, mechanically robust, and require typical operating voltages below 100 V. These properties make SiPMs ideal for replacing PMTs, especially for small-scale space-based detectors. As such a combination of SiPMs and scintillators have been used or proposed for missions such as: GRBAlpha (Pál et al., 2023), VZLUSAT-2 (Münz et al., 2024), BurstCube (Myers et al., 2024), Glowbug (Grove et al., 2020) and GALI (Rahin et al., 2020) which use SiPMs with CsI, COMCUBE (Laviron, 2022) and EIRSAT-1 (Murphy et al., 2021) which are designed to use SiPMs with CeBr3, GARI (Mitchell et al., 2023), GRID (Wen et al., 2019) and GTM (Chang et al., 2022) which use SiPMs coupled to GAGG, MAMBO (Vestrand et al., 2020) which uses it together with BGO and SIRI (Mitchell et al., 2019) which couples it to SrI2. Apart from these relatively small-scale instruments, larger-scale missions like GECAM (Lv et al., 2018), StarBurst (Woolf et al., 2024) and POLAR-2 (Produit et al., 2023b) are also using SiPMs in their design.
@@
In the case of POLAR-2, a total of 6400 scintillators are coupled to an equal number of SiPM channels. This large number of channels required the development of a low-power design. In addition, to keep such a large mission within a manageable budget, as well as to avoid typical export issues that come with high-end space-qualified components, it was decided to only use commercial off-the-shelf (COTS) components. The final design of this front-end readout electronics (FEE), which will be presented in detail in the following section, allows for the readout of 64 SiPM channels. The typical power consumption of the system is 1.8 W, while the cost, excluding SiPMs, is approximately 3000 USD.
@@
While the FEE was designed with the primary goal in mind to be used for POLAR-2, it was soon recognized that it could be used for a wider variety of SiPM-based detectors. For this purpose, the design was kept flexible. For example, while in POLAR-2 it is used to read out a tightly packed plastic scintillator array, it was shown to also function well when reading out different types of scintillators such as GAGG, or CeBr3. While using relatively simple modifications, the layout of the scintillator array, as well as the SiPMs attached can be changed to optimize it for different purposes. The main components of the FEE, as well as the overall design, were furthermore space-qualified during a range of campaigns. As such, the design is currently being proposed to be used on a secondary payload of POLAR-2, the Broad energy-band Spectrum Detector (BSD), an instrument to measure the cosmic X-ray background CXBe (Li et al., 2023; Produit et al., 2023a) and it is being considered for use on the future eXTP mission (Zhang et al., 2019) as a Wide-field Wide energy-band Camera (W2C). Finally, studies to investigate whether the design is suitable for use with linearly graded SiPMs from FBK (LG-SiPMs) (Gola et al., 2020) are currently ongoing by the authors.
@@
In this manuscript, we will present the overall design of the FEE in section 2. This is followed by the details on the performance of the electronics of the first prototypes in section 3. The system was extensively tested to read out Hamamatsu S13661 MPPCs/SiPMs coupled to plastic scintillators, the performance of which is summarized in section 4. The performance after replacing the plastic with GAGG is described in section 5. Finally, the various space qualification tests and their outcomes are presented in section 6. The manuscript finishes with a discussion on potential future applications of the design and shortcomings of the current design.
@@ 2.1 Design constraints
When designed for POLAR-2, the readout system was required to be capable of reading out 64 SiPM channels at low power, low cost and with minimal heating of the SiPMs by the electronics. The low power constraint here stemmed from the power limit of POLAR-2 being 300 W. To optimize the sensitivity of the polarimeter this power needed to be used for as many SiPM channels as possible, the ambitious aim being a total of 6400 channels. The requirement for minimal cost was, as can be expected, related to a tight budget. Finally, the heating requirements stems from SiPMs performing better at lower temperatures. In order to allow for further lowering the temperature of the SiPMs the board is capable of driving a Peltier element which can be placed on the back of the SiPMs. In addition, it can also drive power resistors which have the goal of heating the SiPMs. This capability was added to allow for thermal annealing of radiation damage (De Angelis et al., 2023).
@@
Figure 1: The FEE design in its current form as used for POLAR-2. The two sides are shown here (top and bottom). The leftmost PCB contains the SiPMs, filters for the SiPM bias and heating resistors. This board is connected through a flex to the main PCB which contains the majority of the components. The right PCB contains the connectors as well as drivers which can be used to power a potential Peltier element. (taken from De Angelis (2023), with permission)
@@
Based on the thermal constraints, it was decided to place the main electronics on a PCB separate from the SiPMs, thereby allowing to thermally isolate both parts. The two PCBs are connected through a flexible PCB. Finally, for mechanical convenience, it was then decided to place the various connectors on a final third PCB which again is connected through a flex PCB. The full design can be observed in figure 1, while the way it is mechanically mounted in POLAR-2 is shown in Figure 2.
@@
In order to minimize the power consumption and simplify the design, it was decided to read out the SiPMs using well-tested commercially available ASICs. For this purpose, the Citiroc 1A ASIC by Weeroc (Fleury et al., 2014) was chosen. These 32-channel ASICs have two separate amplifiers for each channel, as well as two separately adjustable thresholds. This offers a large flexibility when designing the trigger logic. More details on the usage of these ASICs is provided in the next section.
@@
Finally, for simplicity, the system is setup to operate on a single voltage which is 3.8 V. The board provides the high voltage at tens of volts (see 2.2.6) for the SiPMs by itself to further reduce dependency on a back-end.
Figure 2: Image of how the front-end electronics is, as of the design in 2023, mechanically integrated in the POLAR-2 detector. The SiPM PCB and connector PCB parts are placed with a 90 angle with respect to the rigid PCB part which holds the main components. This part is mechanically coupled to an aluminum frame in order to thermally insulate it from the SiPMs. In this design the SiPMs are furthermore coupled to a Peltier element connected to a copper bar to remove heat. The SiPMs are connected to a plastic grid which provides precise alignment to the scintillator array (shown in blue at the bottom).
@@ 2.2.2 SiPM PCB
One of the rigid parts contains the SiPMs on one side, while the other side is occupied by filters for the bias voltage, as well as temperature sensors (NTC) and heating resistors. The latter are placed such that the SiPMs can be heated temporarily to speed up the thermal annealing of radiation damage (De Angelis et al., 2023).
In the baseline design, this PCB provides space for 4 SiPM arrays of 4$\times$4 S13661-6075 MPPCs. For POLAR-2 this SiPM type was selected as the detector requires a densely packed readout of plastic scintillator bars. The use of arrays allows minimizing the pitch between scintillators, while the use of the 75 $\mu m$ microcell SiPMs provides the largest possible photo-detection efficiency, typically required for plastic scintillators. As will be discussed in section 5, smaller size microcell SiPMs are recommended when using more luminous scintillator types. This would require no change on the design. If however, the size of the SiPMs (here $6\times 6\,\mathrm{mm^{2}}$, or their pitch needs to be changed, a redesign of the PCB is required. However, as this PCB is basically independent of the significantly more complex part which contains the main electronics, this is relatively straight forward. Only the dimensions need to be modified while the interface with the main PCB can remain the same.
@@ 2.2.3 Connector PCB
The system contains an additional PCB which mainly hosts the various connectors. In the POLAR-2 design, this part additionally hosts drivers for both a Peltier element, which can be used to cool the SiPMs as well as for the heating resistors which can be used to heat these. These drivers are operated directly from the FPGA hosted on the main PCB. For the usage of the Peltier and the heater an additional power supply is required. In the baseline design this is 1V which is the only power source not created directly from the 3.8 V input which powers all the other components on the board. Providing the external 1V power is optional here as it is only used for powering the potential Peltier or heating elements.
Both the 3.8 and the 1V input power are provided to the power connector, which is of the type S2SD 2.00 mm Tiger Eye™ from Samtec. This connector is mechanically robust, while also allowing to couple to power cables from Samtec which do not contain any materials prone to outgassing. It is worth noting that in an earlier prototype, the FTS-105-01-L-DV connector from Samtec was used. Although this performed well during all space qualification tests, it coupled to cables which contained significant amounts of PVC which caused concerns for thermal vacuum tests.
@@ 2.2.4 FPGA
The FPGA was selected based on heritage from the POLAR mission (Produit et al., 2018). The POLAR mission used 25 IGLOO FPGAs from Microsemi, all of which survived 2.5 years in low Earth orbit without any issues  [^1] While POLAR stopped operation after 7 months due to a high voltage issue, the electronics still functioned without issues until the instrument was de-orbited in summer of 2019. Based on this performance, for POLAR-2 a slightly more powerful version, the IGLOO AGL1000V2-FGG256 from Microchip Technology was selected. Apart from the heritage of IGLOO FPGAs from POLAR, this FPGA was selected based on the low power consumption, the large number of I/O’s (294) and its low cost (order of $100\,USD$). The large number of I/O’s is required to deal with the significant number of signals from the two ASICs. Overall, only one pair of I/O’s is not occupied in the baseline design.
Using $35\%$ of the logic gates and $72\%$ of the available memory the FPGA currently performes the following tasks:
@@
  • Optional: Data Compression
  • Optional: Control of the Peltier/Heater Drivers
@@ 2.2.5 ASICs
The ASIC chosen for this readout system is the Citiroc 1A by Weeroc (Fleury et al., 2014). This ASIC was designed to readout a total of 32 SiPMs. It was chosen for this project based on the large flexibility it allows for setting up trigger logics, its capabilities to fine-tune several parameters for each of the 32 detector channels and its Technology Readiness Level of 8.
@@
The ASIC has 2 adjustable amplifiers for each of the 32 channels producing a high gain and a low gain output. The high gain output can be used to measure the energy range with individual photo-electrons, while the low gain can be used to extend the dynamic range into the hundreds of keV or even MeV energy range. Both the high gain and low gain signals are provided in multiplexed outputs. The input signal can be tested against two individual adjustable threshold levels, one called the ’charge threshold’, the other called the ’time threshold’. Whether or not one of the channels has exceeded the charge threshold is provided on one output, while for the time threshold the full trigger pattern is provided in multiplexed form.

### v2

@@ Abstract
keywords:
X-ray 1 , Gamma-Ray 2 , Spectrometer , Astrophysics , Polarimeter , SiPM
journal: Nuclear Instruments and Methods A
[…]
[first]organization=University of New Hampshire, Space Science Center,city=Durham, postcode=03824, state=New Hampshire, country=USA
[…]
[third]organization=INAF-IAPS, ,addressline=via del Fosso del Cavaliere 100, city=Rome, postcode=I-00133, country=Italy
[…]
[second]organization=University of Geneva, DPNC,city=Geneva, postcode=1205, state=Geneva, country=Switzerland
[…]
[4]organization=Geneva Observatory, ISDC, University of Geneva,addressline=16, Chemin d’Ecogia, city=Versoix, postcode=1290, state=Geneva, country=Switzerland
[…]
[5]organization=Max-Planck Institute for Extraterrestrial Physics,addressline=Giessenbachstr. 1, city=Garching, postcode=85748, country=Germany
[…]
[6]organization=Institute of Nuclear Physics Polish Academy of Sciences,city=Krakow, postcode=PL-31342, country=Poland
[…]
[7]organization=National Centre for Nuclear Research,addressline=ul. A. Soltana 7, city=Swierk, postcode=05-400 Otwock, country=Poland
[…]
[8]organization=Key Laboratory of Particle Astrophysics, Institute of High Energy Physics, Chinese Academy of Sciences,city=Beijing, postcode=100049, country=China
@@ 1 Introduction
The joint Gravitational Wave (GW) and Gamma-ray Burst (GRB) detection of GW170817/GRB170817A indicated that gamma-ray detectors are vital in the field of multi-messenger astrophysics [1]. These detectors need to be capable of detecting weak transients and alerting the community with temporal, spectral and localization information. This increase in interest in developing such detectors coincided with the emergence of SiPM technology. Traditional gamma-ray detectors, such as BATSE [2], Fermi-GBM [3], and POLAR [4], were based on scintillators read out by Photo-Multiplier Tubes (PMT). Downsides of PMTs are that they are bulky and mechanically fragile. In addition, such instruments required complex high-voltage power supplies capable of delivering high voltage values typically around kV. The development of SiPMs mitigates all such issues as they are compact, mechanically robust, and require typical operating voltages below 100 V. These properties make SiPMs ideal for replacing PMTs, especially for small-scale space-based detectors. As such, a combination of SiPMs and scintillators have been used or proposed for missions such as: GRBAlpha [5], VZLUSAT-2 [6], BurstCube [7], Glowbug [8] and GALI [9] which use SiPMs with CsI, COMCUBE [10] and EIRSAT-1 [11] which are designed to use SiPMs with CeBr3, GARI [12], GRID [13] and GTM [14] which use SiPMs coupled to GAGG, MAMBO [15] which uses it together with BGO and SIRI [16] which couples it to SrI2. Apart from these relatively small-scale instruments, larger-scale missions like GECAM [17], StarBurst [18] and POLAR-2 [19] are also using SiPMs in their design.
@@
In the case of POLAR-2, a total of 6400 scintillators are coupled to an equal number of SiPM channels. This large number of channels required the development of a low-power electronics design. In addition, to keep such a large mission within a manageable budget, as well as to avoid typical export issues that come with high-end space-qualified components, it was decided to only use commercial off-the-shelf (COTS) components. The final design of this front-end readout electronics (FEE), which will be presented in detail in the following section, allows for the readout of 64 SiPM channels. The typical power consumption of the system is 1.8 W, while the cost, excluding SiPMs, is approximately 3000 USD. One exception to the COTS components is the ASIC used. It should, however, be noted that, at least for POLAR-2, this component is free from export restrictions and commercially available.
@@
While the FEE was designed with the primary goal in mind to be used for POLAR-2, it was soon recognized that it could be used for a wider variety of SiPM-based detectors. For this purpose, the design was kept flexible. For example, while in POLAR-2 it is used to read out a tightly packed plastic scintillator array, it was shown to also function well when reading out different types of scintillators such as GAGG, or CeBr3. This indicates that, with simple modifications, the layout of the scintillar array and the SiPM types attached can be changed relatively easily with respect to the POLAR-2 one to, for example, to produce a design optimized for spectrometry. The main components of the FEE, as well as the overall design, were furthermore space-qualified during a range of campaigns. As such, the design is currently being proposed to be used on a secondary payload of POLAR-2, the Broad energy-band Spectrum Detector (BSD), an instrument to measure the cosmic X-ray background CXBe [20, 21] and it is being considered for use on the future eXTP mission [22] as a Wide-field Wide energy-band Camera (W2C). Finally, studies to investigate whether the design is suitable for use with Linearly Graded SiPMs from FBK (LG-SiPMs) [23] are currently ongoing by the authors.
@@
In this manuscript, we will present the overall design of the FEE in section 2. This is followed by the details on the performance of the electronics of the first prototypes in section 3. The system was extensively tested to read out Hamamatsu S13361 MPPCs/SiPMs coupled to plastic scintillators, the performance of which is summarized in section 4. The performance after replacing the plastic with GAGG is described in section 5. Finally, the various space qualification tests and their outcomes are presented in section 6. The manuscript finishes with a discussion on potential future applications of the design and shortcomings of the current design.
@@ 2.1 Design constraints
When designed for POLAR-2, the readout system was required to be capable of reading out 64 SiPM channels at low power, low cost and with minimal heating of the SiPMs by the electronics. The low power constraint here stemmed from the power limit of POLAR-2 being 300 W. To optimize the sensitivity of the polarimeter this power needed to be used for as many SiPM channels as possible, the ambitious aim being a total of 6400 channels. This would translate to 100 read-out boards to be used. When taking into account the power required for the back-end electronics, voltage conversion efficiencies and heating, this leaves approximately $\approx$ 2 W for each read out board. The requirement for minimal cost was, as can be expected, related to a tight budget. Finally, the heating requirements stems from SiPMs performing better at lower temperatures. For POLAR-2 a target temperature of $0\,\mathrm{C^{o}}$ was set as it provides an acceptable noise level while keeping the SiPM well within the designed operation range. In order to allow for further lowering the temperature of the SiPMs by several degrees and actively stabilizing the temperature of the SiPMs, the board is capable of driving a Peltier element which can be placed on the back of the SiPMs. In addition, it can also drive power resistors which have the goal of heating the SiPMs. This capability was added to allow for thermal annealing of radiation damage [24]. The heating power reserved for both the cooling and heating in POLAR-2 is approximately 1 W. Moreover, it should be noted that, while for POLAR-2 the Peltier element is used for cooling only [^1] A design decision based on limited space on the PCB., one could redesign its control logic and power supply such that it can also heat the SiPMs, thereby removing the need for having the power resistors.
A final consideration in the design, especially important for space missions, is the mass of the design. For POLAR-2 the total allowed mass of the payload is $300\,\mathrm{kg}$ most of which is taken up by the overall mechanics and the mass of the scintillators. The front-end boards discussed here and their mechanics weigh only 170 grams a piece and therefore only make up $17\,\mathrm{kg}$ of the full payload. When used for a Cubesat like mission, this mass is of course a more important issue, where it could be reduced, for example by further minimizing the aluminium mechanics and the copper cooling tube.
@@
Figure 1: The FEE design in its current form as used for POLAR-2. The two sides are shown here (top and bottom). The leftmost PCB contains the SiPMs, filters for the SiPM bias and heating resistors. This board is connected through a flex to the main PCB which contains the majority of the components. The right PCB contains the connectors as well as drivers which can be used to power a potential Peltier element. (taken from [25], with permission)
@@
To reduce heating of the SiPMs by the electronics, it was decided to place these electronics on a PCB separate from the SiPMs, thereby allowing to thermally isolate both parts. The two PCBs are connected through a flexible PCB. Finally, for mechanical convenience, it was then decided to place the various connectors on a final third PCB which again is connected through a flex PCB. The full design can be observed in figure 1, while the way it is mechanically mounted in POLAR-2 is shown in figure 2.
@@
In order to minimize the power consumption and simplify the design, it was decided to read out the SiPMs using well-tested commercially available ASICs. For this purpose, the Citiroc 1A ASIC by Weeroc [26] was chosen. These 32-channel ASICs have two separate amplifiers for each channel, as well as two separately adjustable thresholds. This offers a large flexibility when designing the trigger logic. More details on the usage of these ASICs is provided in the next section.
@@
Finally, for simplicity, the system is setup to operate on a single voltage which is 3.8 V. The board provides the high voltage at tens of volts (see 2.2.7) for the SiPMs by itself to further reduce dependency on a back-end.
Figure 2: Image of how the front-end electronics is, as of the design in 2023, mechanically integrated in the POLAR-2 detector. The SiPM PCB and connector PCB parts are placed with a 90 angle with respect to the rigid PCB part which holds the main components. This main PCB is mechanically coupled to an aluminum frame which forms the start of a thermal path to a radiator. In order to thermally insulate the SiPM PCB from the aluminium the two are separated by a PEEK base. On the other side the SiPMs are connected to a plastic grid which provides precise alignment to the scintillator array (shown in blue at the bottom). In this design the SiPM PCB is furthermore coupled to a Peltier element connected to a copper bar which forms the thermal path to the aluminium.
@@ 2.2.2 SiPM PCB
One of the rigid parts contains the SiPMs on one side, while the other side is occupied by filters for the bias voltage, as well as temperature sensors (NTC) and heating resistors. The latter are placed such that the SiPMs can be heated temporarily to speed up the thermal annealing of radiation damage [24].
In the baseline design, this PCB provides space for 4 SiPM arrays of 4$\times$4 S13361-6075 MPPCs. For POLAR-2 this SiPM type was selected as the detector requires a densely packed readout of plastic scintillator bars. The use of arrays allows minimizing the pitch between scintillators, while the use of the 75 $\mu m$ microcell SiPMs provides the largest possible photo-detection efficiency, typically required for plastic scintillators. As will be discussed in section 5, smaller size microcell SiPMs are recommended when using more luminous scintillator types. This would require no change on the design. If however, the size of the SiPMs (here $6\times 6\,\mathrm{mm^{2}}$), or their pitch needs to be changed, a redesign of the PCB is required. However, as this PCB is basically independent of the significantly more complex part which contains the main electronics, this is relatively straight-forward. Only the dimensions need to be modified while the interface with the main PCB can remain the same.
@@ 2.2.3 Connector PCB
The system contains an additional PCB which mainly hosts the various connectors. In the POLAR-2 design, this part additionally hosts drivers for both a Peltier element and heating resistors which can be used to alter the temperature of the SiPMs. For POLAR-2 current thermal simulations indicate that the system can raise and lower the temperature by aout $5\,\mathrm{C^{o}}$. These drivers are operated directly from the FPGA hosted on the main PCB. For the usage of the Peltier and the heater an additional power supply is required. In the baseline design this is 1V which is the only power source not created directly from the 3.8 V input which powers all the other components on the board. Providing the external 1V power is optional here as it is only used for powering the potential Peltier or heating elements.
Both the 3.8 and the 1V input power are provided to the power connector, which is of the type S2SD 2.00 mm Tiger Eye™ from Samtec  [^2] https://www.samtec.com/products/s2sd. This connector is mechanically robust, while also allowing to couple to power cables from Samtec which do not contain any materials prone to outgassing. It is worth noting that in an earlier prototype, the FTS-105-01-L-DV connector from Samtec was used. Although this performed well during all space qualification tests, it coupled to cables which contained significant amounts of PVC which caused concerns for thermal vacuum tests.
@@ 2.2.4 FPGA
The FPGA was selected based on heritage from the POLAR mission [4]. The POLAR mission used 25 IGLOO FPGAs from Microsemi, all of which survived 2.5 years in low Earth orbit without any issues [^3] While POLAR stopped operation after 7 months due to a high voltage issue, the electronics still functioned without issues until the instrument was de-orbited in summer of 2019.. Based on this performance, for POLAR-2 a slightly more powerful version, the IGLOO AGL1000V2-FGG256 from Microchip Technology was selected. Apart from the heritage of IGLOO FPGAs from POLAR, this FPGA was selected based on the low power consumption, the large number of I/O’s (294) and its low cost (order of $100\,\mathrm{USD}$). The large number of I/O’s is required to deal with the significant number of signals from the two ASICs. Overall, only one pair of I/O’s is not occupied in the baseline design.
Using $35\%$ of the logic gates and $72\%$ of the available memory the FPGA currently performs the following tasks:
  • 1.
@@
  • 11.
    Data Compression
  • 12.
    Control of the Peltier/Heater Drivers
@@ 2.2.5 ASICs
The ASIC chosen for this readout system is the Citiroc 1A by Weeroc [26]. This ASIC was designed to readout a total of 32 SiPMs. It was chosen for this project based on the large flexibility it allows for setting up trigger logics, its capabilities to fine-tune several parameters for each of the 32 detector channels and its Technology Readiness Level of 8.
@@
The ASIC has 2 adjustable amplifiers for each of the 32 channels producing a high gain and a low gain output. Latching of the signal is performed based on comparisons with the various thresholds, or, can be forced externally. The high gain output can be used to measure the energy range with individual photo-electrons, while the low gain can be used to extend the dynamic range into the hundreds of keV or even MeV energy range. Both the high gain and low gain signals are provided in multiplexed outputs. The input signal can be tested against two individual adjustable threshold levels, one called the ’charge threshold’, the other called the ’time threshold’. Although here these two threshold values are used simply as two different levels against which we compare the input signal, and not specifically for charge or time measurements, we will continue to follow this naming convention used by the ASIC manufacturer. If one of the channels has exceeded the charge threshold this information is provided on one output line. In case the time threshold is exceeded by a channel the full trigger pattern is provided in multiplexed form. Details on the complex trigger scheme implemented for the POLAR-2 mission are provided in section 2.2.6.
Apart from its capability to provide a complex, flexible trigger logic, the Citiroc 1A also allows to fine-tune a range of parameters for the 32 channels. These parameters, and whether they are applied to all 32 channels (common), or to individual channels are described in table 1.
Table 1: Your caption.
Parameter Channel/Common Bits
High Gain Amplification Channel 8
Low Gain Amplification Channel 8
Charge Threshold Common 10
Charge Threshold correction Channel 6
Time Threshold Common 10
Time Threshold Correction Channel 6
Bias Voltage Correction Channel 10
The bias voltage correction allows for a common bias voltage can be applied to the 32 channels. The corrections are then used to accommodate for variations in the breakdown between various SiPMs. As such variations, within a single type, are typically of the order of hundreds of mV, the 2.5 V option is normally sufficient to produce a uniform over-voltage. The 4.5 V option can be of use in case, for example, one wants to operate different channels at different over-voltage levels.
It should finally be noted that since the start of the development of this front-end, the Radioroc 2 ASIC has been developed by Weeroc [27]. This ASIC has all the features of the Citiroc 1A while allowing it to read out 64 channels compared to the 32 from the Citiroc. An upgrade to this ASIC in future versions is therefore being considered.
2.2.6 Current Trigger Scheme
The two different amplifiers on the ASIC as well as the two thresholds allow to setup a complex trigger scheme. The system was originally designed for a gamma-ray polarimeter for which coincident events between at least 2 channels, out of the total of 6400, are of interest. These 2 detector channels can be within one 64-channel front-end readout, or can be between 2 different front-ends. To accommodate this, the time threshold and charge threshold are used in parallel to identify different trigger types. In addition, for this experiment, the system is setup to allow to force a trigger externally through an LVDS line which is controlled by a back-end electronics. In all cases, the FPGA forces both ASICs to enter peak sensing mode for 90 ns. During this period each of the 32 channels will record the highest peak, latch on this, and pass this on to the multiplexed outputs (both for HG and LG). The output is passed to ADCs, described later in this section, which convert these pulse heights into digital values before passing these to the FGPA.